OpenStudy (anonymous):

Use the quadratic formula to solve the equation. -2x^2+3x+5=0

hartnn (hartnn):

Compare your quadratic equation with \(ax^2+bx+c=0\) find \[a=...?\\b=...?\\c=...?\\\] \[ \\ \sqrt{b^2-4ac}=...?\] then the two roots of x are: \(\huge{x_{1,2}=\frac{-b \pm \sqrt{b^2-4ac}}{2a}}\)

OpenStudy (anonymous):

somethings not coming up correctly :(. I tried plugging everything in.

hartnn (hartnn):

can you plz show your work?

OpenStudy (anonymous):

\[\sqrt{3x^2-4(2x^2*5)}\]

hartnn (hartnn):

okk...we can't have x in the values of a,b,c